# Correlations for the interphase drag in the two-fluid model of gas-liquid flows through packed-bed reactors

### v2

@@ Abstract
Experiments conducted by NASA measured the pressure drop due to gas-liquid flow through a packed-bed reactor under microgravity conditions. From these experiments, we develop correlations for the gas-liquid $f_{gl}$ interphase drag in a two-fluid model (TFM). We use an Ergun-type closure for liquid-solid drag. Then, under a 1D flow assumption, $f_{gl}$ is the only unknown in the TFM. Using a data-driven approach, we determine $f_{gl}$ and correlate it (via composite fits) with the liquid and gas Reynolds numbers, $Re_{l}$ and $Re_{g}$, respectively, and the Suratman number $Su_{l}$. To validate the proposed $f_{gl}(Re_{l},Re_{g},Su_{l})$ closure, we perform two-dimensional transient simulations at microgravity conditions using ANSYS Fluent and employing an Euler-Euler formulation. We find good agreement between the simulations based on the proposed $f_{gl}$ closure and the experimental data.
@@ 1 Introduction
Multiphase flows through porous media 1, 2 find applications across various systems, from heat pipes to volatile removal assembly and hydrogenation plants, to name a few 3. One such problem is gas-liquid (two-phase) flow through a packed-bed reactor (PBR). A typical PBR consists of spherical catalyst particles packed within a cylindrical (or cuboidal) pipe 4. In the presence of gravity, the gas and liquid phases arrange themselves in the trickle flow regime (the liquid phase trickles over the packing in a gas continuous medium), and the corresponding pressure drops and trickle-to-pulse flow regime transitions have been studied extensively (see, for instance, Refs. 5, 6, 7, 8 and those therein). On the other hand, in the absence of gravity (or under microgravity conditions), the capillary force becomes important. Bubble flow (gas droplets in liquid continuous medium) and pulse flow (alternating regions of gas-rich and liquid-rich) regimes are observed in two-phase flow through PBRs at low-gravity conditions 9, 10, 11. However, determining the pressure drop across the PBR at microgravity conditions remains an active research problem 12. Recently, progress has been made via experimental and modeling efforts.
@@
Specifically, Motil et al. 10 reported on parabolic flight experiments of gas-liquid flow through a PBR. By developing flow regime maps, they demonstrated that different flow physics compete in microgravity conditions compared to terrestrial gravity conditions. In addition, a correlation known as the “modified Ergun equation” was proposed to predict the frictional pressure drop for gas-liquid flows through PBRs at microgravity conditions. The duration of low-gravity conditions on parabolic flights is short. To remedy this shortcoming and gather data on longer-duration PBR flows, further experiments were conducted aboard the International Space Station (ISS)9, 11. Specifically, the wettability and hysteresis effects were studied in the viscous-capillary regime (low Reynolds number flow) by measuring pressure drops and bubble-to-pulse flow regime transitions for Teflon (nonwetting) and glass (wetting) packing materials 9. These experiments suggested that contributions from capillarity dominate the pressure drop for wetting packing, whereas contributions from viscous drag dominate the pressure drop for the nonwetting cases. Further, the effect of pre-flow (i.e., gas-flush versus liquid-flush) through a PBR on the pressure drop was investigated 11. Higher gas holdup and pressure drops were observed for liquid pre-flow conditions compared to gas pre-flow conditions. More recently, a technique to characterize flow regime transitions via the slope of the pressure drop versus flow rate curve (from the ISS data) demonstrated the dominance of capillarity at low flow rates and the dependence of the gas holdup on pre-flow conditions 13.
@@
One-dimensional (1D) two-fluid models (TFMs) have been developed to understand these experimental observations from a fundamental (continuum) point of view 14. Assuming the flow to be 1D, uniform in the cross-section, and at steady-state, the full TFM equations can be simplified so that the unknowns are the axial pressure drop, the liquid holdup, and the interphase drag force 15, 16, 17. Following this approach, Zhang et al. 15 performed experiments and developed a 1D TFM at low Bond number $Bo$ (the ratio of gravitational to surface tension forces), within which the trickle flow regime is dominant. They incorporated Ergun-type correlations in the three interphase drag forces (i.e., the liquid-solid $f_{ls}$, gas-liquid $f_{gl}$, and gas-solid $f_{gs}$ ones) with appropriate relative velocities 6, 7, 8. At low $Bo$, capillarity implies a pressure jump (drop) across a gas-liquid interface: $p_{c}=p_{g}-p_{l}\neq 0$, where $p$ is pressure and the subscripts “$c$”, “$g$”, and “$l$” stand for “capillary,” “gas”, and “liquid,” respectively. For example, Zhang et al. 15 employed the correlation from Attou and Ferschneider 18 to model $p_{c}$ as a function of surface tension, particle diameter, gas holdup, and the ratio of gas and liquid densities. Next, estimating $dp_{c}/dx$ as $p_{c}/d_{p}$, where $p_{c}$ comes from the previous correlation, the respective pressures in the 1D gas and liquid phases’ momentum equations were eliminated to obtain a single momentum equation, with the liquid holdup being the only unknown. Finally, the pressure drop across the PBR was calculated by substituting the predicted liquid holdups, which were determined for different gas and liquid flow rates, obtaining a good agreement between the 1D TFM predictions and the experiments.
@@
In addition to theoretical efforts, computational fluid dynamics (CFD) has been widely used to study resolved two-phase flows through PBRs. Specifically, an Euler-Euler TFM approach is used to predict pressure drop and liquid holdup within PBRs. Previous CFD studies19, 20, 21, 22, 23 considered an “empty” channel and incorporated the relevant closure relations for interphase drag terms (such as $f_{gl}$, $f_{ls}$, and $f_{gs}$) as momentum source terms, which implicitly accounts for the porous medium. In this modeling approach, one does not have to resolve the complex flow domain of interstitial void spaces to extract relevant flow parameters 24. In the present study, we follow this approach and incorporate the proposed $f_{gl}$ closure relations (developed from the 1D TFM model) as momentum source terms in CFD simulation. Then, we compare the pressure drop from these CFD simulations with the experimental pressure drop measured by NASA’s Packed Bed Reactor Experiment (PBRE), which accessed only the bubble and pulse flow regimes at microgravity, provided as data in NASA’s Physical Sciences Informatics (PSI) system 25. This comparison serves as an example demonstration of the utility of the proposed $f_{gl}$ correlations.
To this end, in Sec. 2, we introduce the governing equations (Sec. 2.1) and the data-driven development of correlations for $f_{gl}$ (Sec. 2.2). Then, in Sec. 3, we showcase the utility of the proposed $f_{gl}$ correlation in a CFD simulation (Sec. 3.1) for two different scenarios — wetting and nonwetting packing materials. We compare the pressure drop calculated by CFD against the experimental ones (Sec. 3.2). Finally, Sec. 4 concludes the study.
@@ 2 Mathematical model
Our specific aim is to obtain a closure relation for interphase gas-liquid drag force $f_{gl}$ in a TFM for flow through a PBR, based on the PBRE microgravity experiments 25. The idea of a TFM is to treat the carrier and dispersed phases of a two-phase flow as interpenetrating continua, each having its own thermophysical properties and obeying suitable balance laws 14. The two continua “interact” through the interphase drag force, denoted $f$. Therefore, the interphase drag force couples the momentum equations of the two phases. Throughout, it should be understood that $f$ usually represents a body force density (i.e., a force per unit volume), but we refer to it as “force” for clarity and simplicity.
In this section, we present our proposed procedure for obtaining $f_{gl}$. Specifically, in Sec. 2.1, we start from the basic mass and momentum balances under the TFM. In Sec. 2.2, we employ experimental data from NASA’s PBRE 25 to perform a data-driven inference and calibration of a model for $f_{gl}$. In Sec. 2.3, we demonstrate the self-consistency of the modeling approach by comparing the predicted pressure drops with the original experimental data.
@@ 2.1 Governing equations
The governing equations of gas-liquid flow through a PBR are those of conservation of mass and conservation of momentum. Following the predominant approach in the existing literature 16, 17, 15, and in the absence of detailed measurements of the interstitial flow and interface dynamics, the flow is considered predominantly 1D, with only the gradients along the axial (i.e., flow) direction of PBR being of interest in predicting the pressure drop. Under this assumption, the conservation of mass for liquid and gas phases, respectively, can then be simplified to:
@@
To close the TFM (Eqs. (1) and (2)), we need to specify the interphase momentum transfer terms, which are represented by the forces $f_{g}$ and $f_{l}$. Note that the expressions for $f_{g}$ and $f_{l}$ can each have a number of additive contributions, depending on the physics at hand. To arrive at expressions for $f_{g}$ and $f_{l}$, we assume that the interaction between the gas phase and the solid packing is negligible, i.e., $f_{gs}=0$, because the liquid wets the solid packing 16, 17. Then, we write the phase-specific forces in terms of the interphase ones as $f_{l}=f_{gl}-f_{ls}$ and $f_{g}=-f_{gl}$. We remind the reader that $f_{ls}$ is the interphase drag force between the liquid and solid phases, and $f_{gl}$ is the interphase drag force between the gas and liquid phases.
@@ 2.2 Interphase drag forces
Next, we construct correlations for the interphase drag forces, $f_{gl}$ in Sec. 2.2.1 and $f_{ls}$ in Sec. 2.2.2, required to close Eqs. (3).
2.2.1 Two-phase Ergun correlation for the liquid-solid drag
Previous studies16, 17 used an Ergun-type 26, 27 equation to determine $f_{ls}$. An Ergun-type equation consists of additive viscous and inertial contributions to the pressure gradient across a PBR, but the relation can also be used to express the drag. Specifically, Salgi and Balakotaiah 16, Salgi and Balakotaiah 17 adapted the Ergun-type equation:
@@ 2.2.1 Two-phase Ergun correlation for the liquid-solid drag
2.2.2 Data-driven calibration of the gas-liquid drag
Next, describe our approach to obtaining the closure relation (i.e., correlation) for $f_{gl}$ from the experimental pressure drop data obtained from NASA’s PBRE, which we accessed via NASA’s PSI 25. As indicated in the works of Salgi and Balakotaiah 16, bubble and pulse flow regimes are predominant at microgravity conditions, and the trickle flow regime does not occur. Hence, one cannot use an Ergun-type correlation for $f_{gl}$, with a relative velocity because that is valid only in the trickle flow regime 6, 7, 8. To circumvent this issue, we use a traditional drag force formulation: $f_{gl}=\beta(v_{g}-v_{l})$, where $\beta$ is the drag coefficient, and $v_{g}-v_{l}$ is the relative velocity. Using this formulation, the shared pressure gradient can be eliminated between Eqs. (3a) and (3b), while $f_{ls}$ is eliminated via Eq. (4), to obtain a single governing equation:
@@ 2.2.2 Data-driven calibration of the gas-liquid drag
Salgi and Balakotaiah 16 noted the absence of an acceptable correlation for the capillary pressure $p_{c}$ at microgravity conditions. While previous work 15 estimated $dp_{c}/dx$ as $p_{c}/d_{p}$ using a correlation based on terrestrial gravity conditions for $p_{c}$, we follow the reasoning of Attou and Ferschneider 18, namely “for a fully established gas-liquid flow, the gradient of the capillary pressure can be reasonably neglected compared with the gradient of pressure in each fluid,” justifying the shared pressure gradient assumption that led to Eq. (5).
@@
Using $v_{ls}$ and $v_{gs}$ as inputs and $\phi_{l}$ as output, and along with the PBR’s properties, Eq. (5) has only $\beta$ as an unknown. We use the lsqnonlin function of Matlab to calculate $\beta$ from Eq. (5) for each set of conditions processed from the PBRE datasets. Specifically, we calculate $\beta$ for each set of input velocities. Then, turning the input velocities into Reynolds numbers, we calculate $f_{gl}=\beta(v_{g}-v_{l})$ as a function of the input $Re^{*}_{gs}$ and $Re^{*}_{ls}$ for the Teflon packing dataset. Finally, we plot $f_{gl}$ versus $Re^{*}_{gs}$ in Fig. 1 and observe a linear dependence for each $Re^{*}_{ls}$ (represented by color).
@@
Figure 1: (a) Variation of $f_{gl}$ with $Re^{*}_{gs}$ for the Teflon packing and corresponding linear fits of the form: $f_{gl}(Re^{*}_{ls},Re^{*}_{gs})=a(Re^{*2}_{ls}/Su_{l})+b(Re^{*2}_{ls}/Su_{l})Re^{*}_{gs}$. The variation of $Re^{*}_{ls}$ is represented by color. The arrow indicates the direction of increasing $Re^{*}_{ls}$. (b) Linear fit for the intercepts of $f_{gl}$ of the form: $a(Re^{*2}_{ls}/Su_{l})=\alpha+\beta Re^{*2}_{ls}/Su_{l}$. (c) Linear fit for the slopes of $f_{gl}$ of the form: $b(Re^{*2}_{ls}/Su_{l})=\gamma+\delta Re^{*2}_{ls}/Su_{l}$.
@@
Figure 2: (a) Variation of $f_{gl}$ with $Re^{*}_{gs}$ for the glass packing and corresponding linear fits of the form: $f_{gl}(Re^{*}_{ls},Re^{*}_{gs})=a(Re^{*2}_{ls}/Su_{l})+b(Re^{*2}_{ls}/Su_{l})Re^{*}_{gs}$. The variation of $Re^{*}_{ls}$ is represented by color. The arrow indicates the direction of increasing $Re^{*}_{ls}$. (b) Linear fit for the intercepts of $f_{gl}$ of the form: $a(Re^{*2}_{ls}/Su_{l})=\alpha+\beta Re^{*2}_{ls}/Su_{l}$. (c) Linear fit for the slopes of $f_{gl}$ of the form: $b(Re^{*2}_{ls}/Su_{l})=\gamma+\delta Re^{*2}_{ls}/Su_{l}$.
@@
To quantify this dependence, we performed a set of linear curve fits on the $f_{gl}$ versus $Re^{*}_{gs}$ data at each $Re^{*}_{ls}$ to obtain correlations of the form: $f_{gl}=a+bRe^{*}_{gs}$, where $a$ and $b$ are, respectively, the intercept and slope of the curvefit at fixed value of $Re^{*}_{ls}$. To combine these multiple correlations into a single correlation, we perform a composite linear fit to the coefficients $a$ and $b$ (obtained across all $Re^{*}_{ls}$ values) with respect to $Re^{*2}_{ls}/Su_{l}$, where $Su_{l}=\rho_{l}d_{p}\sigma/\mu_{l}^{2}$ is the Suratman number, as shown in Fig. 1(b) and Fig. 1(c) of the forms: $a(\xi)=\alpha+\beta\xi$ and $b(\xi)=\gamma+\delta\xi$. We introduced $Su_{l}$ (a combined measure of the relative importance of surface tension to both inertial and viscous forces in the flow) here to implicitly account for surface tension (capillarity effects) between the gas and liquid phases, even if the capillary pressure gradient can be neglected in fully developed steady flow 18. Further, following the ideas leading Motil et al. 10 to their dynamic interaction term, we correlate $a$ and $b$ as functions of $Re^{*2}_{ls}/Su_{l}\sim We_{l}$ (the Weber number). Figure 1(a) shows that for largest values of $Re^{*}_{ls}$, the dependence of $f_{gl}$ on $Re^{*}_{gs}$ tends to be nonlinear for small values of $Re^{*}_{gs}$. However, since we perform a curve fit over the full range of data, these outliers do not significantly affect the trend and are not handled differently (see also the discussion of Fig. 3 below). In principle, fitting functions with more parameters could be devised to capture those points better, but given the uncertainties involved at low flow rates, such an approach would likely result in overfitting.
@@
Finally, if one were to justify that the capillary pressure gradient in a two-phase flow in microgravity conditions is not negligible, it can be added back into Eq. (5), potentially via a correlation of the form $p_{c}/d_{p}$ 15. In such a case, the fitting process can be repeated, and the curve-fit coefficients would change but the methodology remains the same. It must also be noted that the proposed correlation for $f_{gl}$ is applicable for monodisperse Teflon and glass packing materials and cannot be generalized for other packing materials with different sizes. This drawback stems from the limited nature of the experimental data from NASA’s PBRE 25. Furthermore, the proposed correlations can be used to predict bulk properties, such as the gas-liquid pressure drop across a PBR at microgravity conditions, but not interstitial properties, such as the flow regime, which are limited to bubble and pulse flow in NASA’s PBRE at microgravity conditions.
Figure 3: Verification of the proposed $f_{gl}$ correlations in Eq. (6) (and hence the approach) by comparing the model and experiment’s $-dp/dz$ variation with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon packing and (b) glass packing PBRE data. As expected, the model predictions ($\bigtriangleup$) at discrete data points coincide exactly with the values from experiments ($\times$). The arrow indicates the direction of increasing $Re^{*}_{ls}$. The dashed lines are a guide to the eye and do not represent any trends. Note the logarithmic abscissa.

### v3

# Correlations for the Interphase Drag in the Two-Fluid Model of Gas–Liquid Flows through Packed-Bed Reactors

@@ Abstract
Experiments conducted by NASA measured the pressure drop due to gas–liquid flow through a packed-bed reactor under microgravity conditions. From these experiments, we develop correlations for the gas–liquid $f_{gl}$ interphase drag in a two-fluid model (TFM). We use an Ergun-type closure for liquid–solid drag. Then, under a 1D flow assumption, $f_{gl}$ is the only unknown in the TFM. Using a data-driven approach, we determine $f_{gl}$ and correlate it (via composite fits) with the liquid and gas Reynolds numbers, $Re_{l}$ and $Re_{g}$, respectively, and the Suratman number $Su_{l}$. To validate the proposed $f_{gl}(Re_{l},Re_{g},Su_{l})$ closure, we perform two-dimensional transient simulations at microgravity conditions using ANSYS Fluent and employing an Euler–Euler formulation. We find good agreement between the simulations based on the proposed $f_{gl}$ closure and the experimental data.
@@ 1 Introduction
Multiphase flows through porous media 1, 2 find applications across various systems, from heat pipes to volatile removal assembly and hydrogenation plants, to name a few 3. One such problem is gas–liquid (two-phase) flow through a packed-bed reactor (PBR). A typical PBR consists of spherical catalyst particles packed within a cylindrical (or cuboidal) pipe 4. In the presence of gravity, the gas and liquid phases arrange themselves in the trickle flow regime (the liquid phase trickles over the packing in a gas continuous medium), and the corresponding pressure drops and trickle-to-pulse flow regime transitions have been studied extensively (see, for instance, Refs. 5, 6, 7, 8 and those therein). On the other hand, in the absence of gravity (or under microgravity conditions), the capillary force becomes important. Bubble flow (gas droplets in liquid continuous medium) and pulse flow (alternating regions of gas-rich and liquid-rich) regimes are observed in two-phase flow through PBRs at low-gravity conditions 9, 10, 11. However, determining the pressure drop across the PBR at microgravity conditions remains an active research problem 12. Recently, progress has been made via experimental and modeling efforts.
@@
Specifically, Motil et al. 10 reported on parabolic flight experiments of gas–liquid flow through a PBR. By developing flow regime maps, they demonstrated that different flow physics compete in microgravity conditions compared to terrestrial gravity conditions. In addition, a correlation known as the “modified Ergun equation” was proposed to predict the frictional pressure drop for gas–liquid flows through PBRs at microgravity conditions. The duration of low-gravity conditions on parabolic flights is short. To remedy this shortcoming and gather data on longer-duration PBR flows, further experiments were conducted aboard the International Space Station (ISS)9, 11. Specifically, the wettability and hysteresis effects were studied in the viscous-capillary regime (low Reynolds number flow) by measuring pressure drops and bubble-to-pulse flow regime transitions for Teflon (nonwetting) and glass (wetting) packing materials 9. These experiments suggested that contributions from capillarity dominate the pressure drop for wetting packing, whereas contributions from viscous drag dominate the pressure drop for the nonwetting cases. Further, the effect of preflow (i.e., gas-flush versus liquid-flush) through a PBR on the pressure drop was investigated 11. Higher gas holdup and pressure drops were observed for liquid preflow conditions compared to gas preflow conditions. More recently, a technique to characterize flow regime transitions via the slope of the pressure drop versus flow rate curve (from the ISS data) demonstrated the dominance of capillarity at low flow rates and the dependence of the gas holdup on preflow conditions 13.
@@
One-dimensional (1D) two-fluid models (TFMs) have been developed to understand these experimental observations from a fundamental (continuum) point of view 14. Assuming the flow to be 1D, uniform in the cross-section, and at steady-state, the full TFM equations can be simplified so that the unknowns are the axial pressure drop, the liquid holdup, and the interphase drag force 15, 16, 17. Following this approach, Zhang et al. 15 performed experiments and developed a 1D TFM at low Bond number $Bo$ (the ratio of gravitational to surface tension forces), within which the trickle flow regime is dominant. They incorporated Ergun-type correlations in the three interphase drag forces (i.e., the liquid–solid $f_{ls}$, gas–liquid $f_{gl}$, and gas–solid $f_{gs}$ ones) with appropriate relative velocities 6, 7, 8. At low $Bo$, capillarity implies a pressure jump (drop) across a gas–liquid interface: $p_{c}=p_{g}-p_{l}\neq 0$, where $p$ is pressure and the subscripts “$c$”, “$g$”, and “$l$” stand for “capillary,” “gas”, and “liquid,” respectively. For example, Zhang et al. 15 employed the correlation from Attou and Ferschneider 18 to model $p_{c}$ as a function of surface tension, particle diameter $d_{p}$, gas holdup, and the ratio of gas and liquid densities. Next, estimating $dp_{c}/dz$ as $p_{c}/d_{p}$, where $p_{c}$ comes from the previous correlation, the respective pressures in the 1D gas and liquid phases’ momentum equations were eliminated to obtain a single momentum equation, with the liquid holdup being the only unknown. Finally, the pressure drop across the PBR was calculated by substituting the predicted liquid holdups, which were determined for different gas and liquid flow rates, obtaining a good agreement between the 1D TFM predictions and the experiments.
@@
In addition to theoretical efforts, computational fluid dynamics (CFD) has been widely used to study resolved two-phase flows through PBRs. Specifically, an Euler–Euler TFM approach is used to predict pressure drop and liquid holdup within PBRs. Previous CFD studies19, 20, 21, 22, 23 considered an “empty” channel and incorporated the relevant closure relations for interphase drag terms (such as $f_{gl}$, $f_{ls}$, and $f_{gs}$) as momentum source terms, which implicitly accounts for the porous medium. In this modeling approach, one does not have to resolve the complex flow domain of interstitial void spaces to extract relevant flow parameters 24. In the present study, we follow this approach and incorporate the proposed $f_{gl}$ closure relations (developed from the 1D TFM model) as momentum source terms in CFD simulation. Then, we compare the pressure drop from these CFD simulations with the experimental pressure drop measured by NASA’s Packed Bed Reactor Experiment (PBRE), which accessed only the bubble and pulse flow regimes at microgravity, provided as data in NASA’s Physical Sciences Informatics (PSI) system 25. The PBRE data in the PSI repository 25 consists of two datasets: one for a glass packing (wetting material) and one for a Teflon packing (nonwetting material). For each dataset, the gas and liquid flow rates were fixed (controlled), and the pressure gradient across the PBR was measured. The comparison between the CFD simulations and PBRE data serves as an example demonstration of the utility of the proposed $f_{gl}$ correlations.
This paper is organized as follows. In Sec. 2, we introduce the governing equations (Sec. 2.1) and modeling approach (Sec. 2.2). In Sec. 3, we discuss the data-driven development of correlations for $f_{gl}$. Then, in Sec. 4, we showcase the utility of the proposed $f_{gl}$ correlation in a CFD simulation (Sec. 4.1) for two different scenarios — wetting and nonwetting packing materials, and we compare the pressure drop calculated by CFD against the experimental ones (Sec. 4.2). Section 5 concludes the study.
@@ 2 Mathematical Model
Our specific aim is to obtain a closure relation for interphase gas–liquid drag force $f_{gl}$ in a TFM for flow through a PBR, based on the PBRE microgravity experiments 25. The idea of a TFM is to treat the carrier and dispersed phases of a two-phase flow as interpenetrating continua, each having its own thermophysical properties and obeying suitable balance laws 14. The two continua “interact” through the interphase drag force, denoted $f$. Therefore, the interphase drag force couples the momentum equations of the two phases. Throughout, it should be understood that $f$ usually represents a body force density (i.e., a force per unit volume), but we refer to it as “force” for clarity and simplicity.
In this section, we present our proposed procedure for obtaining $f_{gl}$. Specifically, in Sec. 2.1, we start from the basic mass and momentum balances under the TFM. In Sec. 2.2, we employ experimental data from NASA’s PBRE 25 to perform a data-driven inference and calibration of a model for $f_{gl}$.
@@ 2.1 Governing Equations
The governing equations of gas–liquid flow through a PBR are those of conservation of mass and conservation of momentum. Following the predominant approach in the existing literature 16, 17, 15, and in the absence of detailed measurements of the interstitial flow and interface dynamics, the flow is considered predominantly 1D, with only the gradients along the axial (i.e., flow) direction of PBR being of interest in predicting the pressure drop. Under this assumption, the conservation of mass for liquid and gas phases, respectively, can then be simplified to:
@@
To close the TFM (Eqs. (1) and (2)), we need to specify the interphase momentum transfer terms, which are represented by the forces $f_{g}$ and $f_{l}$. Note that the expressions for $f_{g}$ and $f_{l}$ can each have a number of additive contributions, depending on the physics at hand (such as virtual mass, lift, etc.). In the present work, we focus on the main contribution, i.e., the drag. To arrive at expressions for $f_{g}$ and $f_{l}$, we assume that the interaction between the gas phase and the solid packing is negligible, i.e., $f_{gs}=0$, because the liquid wets the solid packing 16, 17. Then, we write the phase-specific forces in terms of the interphase ones as $f_{l}=f_{gl}-f_{ls}$ and $f_{g}=-f_{gl}$. We remind the reader that $f_{ls}$ is the interphase drag force between the liquid and solid phases, and $f_{gl}$ is the interphase drag force between the gas and liquid phases.
@@ 2.2 Interphase Drag Forces
Next, we construct correlations for the interphase drag forces, $f_{ls}$ in Sec. 2.2.1 and $f_{gl}$ in Sec. 2.2.2, required to close Eqs. (3).
2.2.1 Two-Phase Ergun Correlation for the Liquid–Solid Drag
Previous studies16, 17 used an Ergun-type 26, 27 equation to determine $f_{ls}$. An Ergun-type equation consists of additive viscous and inertial contributions to the pressure gradient across a PBR, but the relation can also be used to express the drag. Specifically, Salgi and Balakotaiah 16, Salgi and Balakotaiah 17 adopted the Ergun-type equation:
@@ 2.2.1 Two-Phase Ergun Correlation for the Liquid–Solid Drag
2.2.2 Formulation for the Gas–Liquid Drag
Next, we describe our approach to obtaining the closure relation (i.e., correlation) for $f_{gl}$ from the experimental pressure drop data obtained from NASA’s PBRE, which we accessed via NASA’s PSI 25. As indicated in the works of Salgi and Balakotaiah 16, bubble and pulse flow regimes are predominant at microgravity conditions, and the trickle flow regime does not occur. Hence, one cannot use an Ergun-type correlation for $f_{gl}$, with a relative velocity because that is valid only in the trickle flow regime 6, 7, 8. To circumvent this issue, we use a traditional drag force formulation: $f_{gl}=\beta(v_{g}-v_{l})$, where $\beta$ is the drag coefficient, and $v_{g}-v_{l}$ is the relative velocity. Using this formulation, the shared pressure gradient can be eliminated between Eqs. (3a) and (3b), while $f_{ls}$ is eliminated via Eq. (4), to obtain a single governing equation:
@@ 2.2.2 Formulation for the Gas–Liquid Drag
Salgi and Balakotaiah 16 noted the absence of an acceptable correlation for the capillary pressure $p_{c}$ at microgravity conditions. While previous work 15 estimated $dp_{c}/dz$ as $p_{c}/d_{p}$ using a correlation based on terrestrial gravity conditions for $p_{c}$, we follow the reasoning of Attou and Ferschneider 18, namely “for a fully established gas–liquid flow, the gradient of the capillary pressure can be reasonably neglected compared with the gradient of pressure in each fluid,” justifying the shared pressure gradient assumption that led to Eq. (5).
3 Data-Driven Calibration of the Mathematical Model
3.1 Description of the PBRE Dataset
The PBRE data in the PSI repository 25 consists of two sets, one for glass packing (wetting) and one for Teflon packing (nonwetting). Each dataset has gas and liquid flow rates as input and the resulting pressure gradient across the PBR as output. The absolute pressure was also measured at five axial locations along the PBR. In the present analysis, we exclude the data in the PBRE data that was explicitly labeled as having significant experimental errors, such as problems with the setup and measurements done aboard the ISS for these cases. We repeat our analysis shown below twice for each of the packing materials used in the PBRE to implicitly capture the contact angle dependence (effect of wetting) in our model.
@@ 3.2 Calibration Procedure
Using $v_{ls}$ and $v_{gs}$ as inputs and $\phi_{l}$ as output, and along with the PBR’s properties, Eq. (5) has only $\beta$ as an unknown. We used the lsqnonlin function of Matlab to calculate $\beta$ from Eq. (5) for each set of conditions processed from the PBRE datasets. Specifically, we calculated $\beta$ for each set of input velocities. Then, turning the input velocities into Reynolds numbers, we calculated $f_{gl}=\beta(v_{g}-v_{l})$ as a function of the input $Re^{*}_{gs}$ and $Re^{*}_{ls}$ for the Teflon packing dataset. Finally, we plot $f_{gl}$ versus $Re^{*}_{gs}$ in Fig. 1 and observe a linear dependence for each $Re^{*}_{ls}$ (represented by color).
@@
Figure 1: (a) Variation of $f_{gl}$ with $Re^{*}_{gs}$ for the Teflon packing and corresponding linear fits of the form: $f_{gl}(Re^{*}_{ls},Re^{*}_{gs})=a(Re^{*2}_{ls}/Su_{l})+b(Re^{*2}_{ls}/Su_{l})Re^{*}_{gs}$. The values of $Re^{*}_{ls}$ (represent by different colors) used are $3.5$, $8.4$, $20.9$, $55.7$, $69.6$, $87.0$, and $104.4$. The arrow represents the direction of increasing $Re^{*}_{ls}$. (b) Linear fit for the intercepts of $f_{gl}$ of the form: $a(Re^{*2}_{ls}/Su_{l})=\alpha+\beta Re^{*2}_{ls}/Su_{l}$. (c) Linear fit for the slopes of $f_{gl}$ of the form: $b(Re^{*2}_{ls}/Su_{l})=\gamma+\delta Re^{*2}_{ls}/Su_{l}$.
@@
Figure 2: (a) Variation of $f_{gl}$ with $Re^{*}_{gs}$ for the glass packing and corresponding linear fits of the form: $f_{gl}(Re^{*}_{ls},Re^{*}_{gs})=a(Re^{*2}_{ls}/Su_{l})+b(Re^{*2}_{ls}/Su_{l})Re^{*}_{gs}$. The values of $Re^{*}_{ls}$ (represent by different colors) used are $3.5$, $8.4$, $20.9$, $55.7$, $69.6$, $87.0$, and $104.4$. The arrow represents the direction of increasing $Re^{*}_{ls}$. (b) Linear fit for the intercepts of $f_{gl}$ of the form: $a(Re^{*2}_{ls}/Su_{l})=\alpha+\beta Re^{*2}_{ls}/Su_{l}$. (c) Linear fit for the slopes of $f_{gl}$ of the form: $b(Re^{*2}_{ls}/Su_{l})=\gamma+\delta Re^{*2}_{ls}/Su_{l}$.
@@
To quantify this dependence, we performed a set of linear curve fits on the $f_{gl}$ versus $Re^{*}_{gs}$ data at each $Re^{*}_{ls}$ to obtain correlations of the form: $f_{gl}=a+bRe^{*}_{gs}$, where $a$ and $b$ are, respectively, the intercept and slope of the curvefit at fixed value of $Re^{*}_{ls}$. To combine these multiple correlations into a single correlation, we performed a composite linear fit to the coefficients $a$ and $b$ (obtained across all $Re^{*}_{ls}$ values) with respect to $Re^{*2}_{ls}/Su_{l}$, where $Su_{l}=\rho_{l}d_{p}\sigma/\mu_{l}^{2}$ is the Suratman number, as shown in Fig. 1(b) and Fig. 1(c) of the forms: $a(\xi)=\alpha+\beta\xi$ and $b(\xi)=\gamma+\delta\xi$. We introduced $Su_{l}$ (a combined measure of the relative importance of surface tension to both inertial and viscous forces in the flow) here to implicitly account for surface tension (capillarity effects) between the gas and liquid phases, even if the capillary pressure gradient can be neglected in fully developed steady flow 18. Further, following the ideas leading Motil et al. 10 to their dynamic interaction term, we correlate $a$ and $b$ as functions of $Re^{*2}_{ls}/Su_{l}\sim We_{l}$ (the Weber number). Figure 1(a) shows that for largest values of $Re^{*}_{ls}$, the dependence of $f_{gl}$ on $Re^{*}_{gs}$ tends to be nonlinear for small values of $Re^{*}_{gs}$. However, since we performed a curve fit over the full range of data, these outliers do not significantly affect the trend and are not handled differently (see also the discussion of Fig. 3 below). In principle, fitting functions with more parameters could be devised to capture those points better, but given the uncertainties involved at low flow rates, such an approach would likely result in overfitting.
@@
Finally, if one were to justify that the capillary pressure gradient in a two-phase flow in microgravity conditions is not negligible, it can be added back into Eq. (5), potentially via a correlation of the form $p_{c}/d_{p}$ 15. In such a case, the fitting process can be repeated, and the curve-fit coefficients would change, but the methodology remains the same. It must also be noted that the proposed correlation for $f_{gl}$ is applicable for monodisperse Teflon and glass packing materials and cannot be generalized for other packing materials with different sizes. This drawback stems from the limited nature of the experimental data from NASA’s PBRE 25. Furthermore, the proposed correlations can be used to predict bulk properties, such as the gas–liquid pressure drop across a PBR at microgravity conditions, but not interstitial properties, such as the flow regime, which are limited to bubble and pulse flow in NASA’s PBRE at microgravity conditions.
Figure 3: Verification of the proposed $f_{gl}$ correlations in Eq. (6) (and hence the approach) by comparing the model and experiment’s $-dp/dz$ variation with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon packing and (b) glass packing PBRE data. As expected, many of the model predictions ($\bigtriangleup$) at discrete data points coincide exactly with the values from experiments ($\times$). The arrow indicates the direction of increasing $Re^{*}_{ls}$. The dashed lines are a guide to the eye and do not represent any trends. Note the logarithmic abscissa.